Circular No. 7031

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET P/1998 S1 On Oct. 17, Jean Mueller reported her discovery of a cometary object

with a strong condensation and a faint but obvious tail to the southwest on a 30-min exposure on Oct. 14 with the 1.2-m Oschin Schmidt telescope at Palomar. The plate was taken by K. Rykoski and Mueller in the course of the Palomar Outer Solar System Ecliptic Survey, with M. Brown as the principal investigator. The comet was confirmed on Oct. 17 on CCD frames obtained by R. Gal with the 1.5-m Palomar telescope and on another Oschin Schmidt 30-min exposure (in seeing so bad that it compromised the positional measurements) by Mueller. G. V. Williams, Minor Planet Center, reduced the Mueller data and found that the comet was probably identical with an apparently quite ordinary minor planet observed by LINEAR on Sept. 26 and 27. He placed information about the object in The NEO Confirmation Page. Initial confirmation of the identity of the LINEAR and Mueller objects was then provided by new measurements made at the Kleť Observatory by J. Tichá and M. Tichý, who remarked that the object was slightly diffuse, with a coma of diameter not less than 13′′ and a faint tail extending for not less than 30′′ in p.a. 210^{o}. His resulting orbit improvement allowed Williams also to identify a single-night (Sept. 17) prediscovery detection in “asteroidal” data from the Lowell Observatory LONEOS program. The comet is of short period and had a close approach to Jupiter in 1992. The discovery positions, orbital elements and ephemeris below are extracted from the more complete data given on MPEC 1998-U07.

1998 UT α_{2000} δ_{2000} Mag. Observer
Sept.26.21764 1^{h}13^{m}33.^{s}81 +12^{o}41′12′′.7 18.1 LINEAR
Oct. 14.25694 1 00 13.71 +13 02 14.8 17  Mueller
T = 1998 Nov. 3.180 TT  ω = 26.^{o}489
e = 0.41690  Ω = 359.183 }2000.0
q = 2.54819 AU  i = 10.559
a = 4.37008 AU n^{o} = 0.107887 P = 9.14 years
1998TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Oct. 14 1^{h}00.^{m}42 +13^{o}02.^{′}2 1.558 2.552 173.^{o}7 2.^{o}4 15.5
24 0 53.06 +13 04.1 1.575 2.549 165.3  5.7 15.6
Nov. 3 0 47.01 +13 04.3 1.617 2.548 154.5  9.6 15.6
13 0 43.03 +13 07.4 1.682 2.549 143.9 13.2 15.7
23 0 41.56 +13 16.9 1.768 2.552 133.7 16.2 15.8
1998 October 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Brian G. Marsden